# POWER-DOMINANCE IN ESTIMATION THEORY: A THIRD PATHOLOGICAL AXIS

### v1

@@ 2.2 Power-dominance Penalty
Let $X,\hat{X}\in L^{2}(\Omega)$ and $e\triangleq\hat{X}-X$. If the estimator operates in the power-dominance regime and $\mathbb{E}[\hat{X}e]\neq 0$, then
@@
with $A=\hat{X}$ and $B=X$, we obtain
$\langle\hat{X},e\rangle$ $=\tfrac{1}{2}\,\mathbb{E}[e^{2}]+\tfrac{1}{2}\big{(}\mathbb{E}[\hat{X}^{2}]-\mathbb{E}[X^{2}]\big{)}.$ (4)
If the estimator is in Regime 1, then from (4) and [6], we get
$\mathbb{E}[\hat{X}\cdot e]>\tfrac{1}{2}\,\mathbb{E}[e^{2}]=\tfrac{1}{2}\,\mathrm{MSE}.$ (5)
@@
Under the same setup as Theorem 2.1, if $\mathbb{E}[\hat{X}^{2}]\leq\mathbb{E}[X^{2}]$ (Regime 2 or Regime 3), then
@@
The Theorem 2.1 and Corollary 2.1.1 establish that the mean power of the estimate relative to the true value is a critical diagnostic for the structural validity of an estimator. This shifts the focus from the classical bias-variance trade-off towards the ‘bias-variance-power triad’ by adding a deeper criterion: The mean power of the estimate must not exceed that of the signal if orthogonality and low-error performance are to coexist. This unifying perspective links estimation theory, information theory, and optimization. It has the potential to inspire new regularization strategies and estimator designs.
@@ 6 Conclusion
This paper introduces a new framework for estimator design based on a second-order diagnostic: the mean power of the estimate relative to the true signal. We classify estimators into three power regimes and show that power-dominant estimators are structurally prone to an unavoidable MSE penalty. This finding offers a new perspective on estimator design, as traditional bias and variance metrics may not fully reveal an estimator’s health. The proposed framework is visually represented by two safe-zone maps. The first provides a diagnostic similar to an ROC curve, while the second shows that the “safe-zone” corresponds to a bounded, well-posed optimization problem, in stark contrast to the unbounded landscape of the “forbidden zone” Our work thus reframes estimator design as a path optimization problem in resource allocation and dynamic power control, providing new avenues for research in regularization and adaptive estimation.

### v2

@@ 2.2 Power-dominance Penalty
Let $X,\hat{X}\in L^{2}(\Omega)$ and $e\triangleq\hat{X}-X$. If the estimator operates in the power-dominance regime and $\mathbb{E}[\hat{X}\cdot e]\neq 0$, then
@@
with $A=\hat{X}$, and $B=X$ as elements in the Hilbert space $L^{2}(\Omega)$, where $\langle.,.\rangle$ corresponds to expectation [6], we get,
$\langle\hat{X},e\rangle$ $=\tfrac{1}{2}\,\mathrm{MSE}+\tfrac{1}{2}\big(\mathbb{E}[\hat{X}^{2}]-\mathbb{E}[X^{2}]\big).$ (4)
In Regime 1, with $\mathbb{E}[\hat{X}^{2}]>\mathbb{E}[X^{2}]$, from (4) and [7], under non-degeneracy, we get, $\mathbb{E}[\hat{X}\cdot e]>\tfrac{1}{2}\,\mathrm{MSE}$. The degeneracy condition $\mathbb{E}[\hat{X}\cdot e]\neq 0$ excludes three cases: (i) trivial $\hat{X}=0$, (ii) optimal $\mathbb{E}[\hat{X}\cdot e]=0$, and (iii) ideal $\hat{X}=X$. ∎
@@
Under the same setup as Theorem 2.1, if $\mathbb{E}[\hat{X}^{2}]\leq\mathbb{E}[X^{2}]$ (Power-conservative regime), then
@@
The Theorem 2.1 and Corollary 2.1.1 establish that the mean power of the estimate relative to the true value is a critical diagnostic for the structural validity of an estimator. This shifts the focus from the classical bias-variance trade-off towards the ‘bias-variance-power triad’ by adding a deeper criterion: The mean power of the estimate must not exceed that of the signal if orthogonality and low-error performance are to coexist.
@@ 6 Conclusion
This paper introduces a new framework for estimator design based on a second-order diagnostic: the mean power of the estimate relative to the true signal. We classify estimators into three power regimes and show that power-dominant estimators are structurally prone to an unavoidable MSE penalty. This finding offers a new perspective on estimator design, as traditional bias and variance metrics may not fully reveal an estimator’s health. This result extends the classical bias–variance trade-off into a ‘bias–variance–power triad,’ positioning mean power as a structural diagnostic for estimator validity. The proposed framework is visually represented by two safe-zone maps. The first provides a diagnostic similar to an ROC curve, while the second shows that the safe-zone” corresponds to a bounded, well-posed optimization problem, in stark contrast to the unbounded landscape of the forbidden zone.” Our work thus reframes estimator design as a path optimization problem in resource allocation and dynamic power control, providing new avenues for research in regularization and adaptive estimation.